OpenStudy (anonymous):

(2^3)3 write in a whole number (2^3)3 write in a whole number @Mathematics

OpenStudy (anonymous):

2^6

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Is it (2^3)^3 or (2^3)3

OpenStudy (anonymous):

Is the three on the outside of the parenthesis an exponent or constant?

OpenStudy (anonymous):

\[(2^3)^3=2^{3\times 3}=2^9=512\]

OpenStudy (anonymous):

(2^3)3 = 24

OpenStudy (anonymous):

LOL wow I fail, yeah its not 12 ^ he is right

OpenStudy (anonymous):

Because if it is(2^3)^3 then the answer is 512 or if it is (2^3)3 the answer to this is 24

OpenStudy (anonymous):

thanks